OpenStudy (anonymous):

I'm literally about to cry from this problem, medal and fan will be rewarded for help. :( Linda is studying the sale of a particular brand of cereals from the year 1993 to 2004. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t) = t^2 + 5t + 52 What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work.

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

First you'll need to calculate S(1) and S(4)

OpenStudy (anonymous):

Right

jimthompson5910 (jim_thompson5910):

what do you get for each?

OpenStudy (anonymous):

I got it, let me type it up:

OpenStudy (anonymous):

\[\frac{ y^2 - y^1 }{ x^2 - x^1 }\]

jimthompson5910 (jim_thompson5910):

right idea, but the notation is off

OpenStudy (anonymous):

The notation? Would you mind explaining a little more? :)

jimthompson5910 (jim_thompson5910):

basically you're calculating the slope through (x1,y1) and (x2,y2) where y1 = S(x1) and y2 = S(x2)

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

well you're not squaring y, it's y sub 2 (the 2 is below the y)

OpenStudy (anonymous):

So, 1997 = S(4), 1993 = S(1)

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

you plug t = 1 into S(t) to find S(1)

OpenStudy (anonymous):

OOOOooooohhhh!

OpenStudy (anonymous):

So:

OpenStudy (anonymous):

uuuuuhhhh.....could you please explain more please?

jimthompson5910 (jim_thompson5910):

S(t) = t^2 + 5t + 52 S(1) = (1)^2 + 5(1) + 52 ... replace EVERY copy of t with 1 S(1) = ???

OpenStudy (anonymous):

Oh poop, that's what I was going to put, but I was thinking of a different equation, hold on one sec, let me solve it! XD

OpenStudy (anonymous):

\[S(1) = 1^2 + 5(1) + 52 = S(1) = 58?\]

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

S(4) = ??

OpenStudy (anonymous):

Yeeeesssss...hehehe....let me solve that one...

OpenStudy (anonymous):

\[S(4) = 4^2 + 5(4) + 52 = S(4) = 88\]

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

OOOOOOOHHHHHHH.....I GOT IT. PLUG THOSE INTO THE AVERAGE RATE OF CHANGE RIGHT?

OpenStudy (anonymous):

The formula?

jimthompson5910 (jim_thompson5910):

exactly, visually the average rate of change is just the slope of the line through the two points

OpenStudy (anonymous):

I understand now! I forgot the AROC is the slope! Ok, plug those into the parts in the formula and I'll get my answer, correct good sir?

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

Alright, I'll answer it, hold on a sec.:

OpenStudy (anonymous):

\[\frac{ 88 - 58 }{ 4 - 1 } = \frac{ 30 }{ 3 } = 10 million dollars!!!!!!\]

jimthompson5910 (jim_thompson5910):

10 million is correct, but the units are dollars per yer

OpenStudy (anonymous):

Mr. Thompson, would you mind deleting your answer about the "now use them to find the slope through (1,58) and (4,88) ?" I don't want other people to see how you got the answer, like just taking the answer without knowing what it means.

jimthompson5910 (jim_thompson5910):

fair enough

OpenStudy (mathmale):

Actually, I think others would benefit from seeing what jim wrote to you. He discusses a method, but does not provide an answer!

jimthompson5910 (jim_thompson5910):

probably best to delete the whole question then (or at least that final answer) if you're concerned about that

OpenStudy (anonymous):

Ok, I shall do that. @mathmale you are correct about that, but I know some students who just look at these questions, and take the answers. :(

OpenStudy (anonymous):

Why? You will help other people?? :\

OpenStudy (anonymous):

I know, but...I guess you're right.

OpenStudy (anonymous):

I won't say the answer. How about that? :)

OpenStudy (mathmale):

Your discussion with Jim has been a very good one. I see no harm done by leaving it posted as is.

OpenStudy (anonymous):

Ok, I hope I didn't break any rules @mathmale

OpenStudy (mathmale):

Relax. Not at all! On the contrary, you were a thoughtful participant in this discussion. What I'd give to see others participate as much as you have! Why not go on to some new problems now? Good luck!

OpenStudy (anonymous):

Taw, thank you for the kind words! I shall continue on my journey for new math problems to conquer. Farewell!